As ASAR was a very successful mission, there were many highlights. Many of these highlights were presented during the ESA Living Planet Symposiums such as the one held in Edinburgh, UK in 2013, the year after contact was lost with ENVISAT. The programme for this symposium, including abstracts, can be found here. Full details of ASAR operational activities throughout the mission are available in the paper from this conference by Miranda et al. (2013): The Envisat ASAR Mission: A look back at 10 years of operation, which is available here.
Standard ASAR operations provided a 35 day repeat orbit. Major events during operations included:
ASAR Data Subsystem Redundancy Switch (May 2003)
First signs of an on-board anomaly were found with the quality of ASAR data just after launch, although no problems were detected from the instrument telemetry. A failure was then found on the Q-branch of the on-board transmitter. ASAR was switched of on 14 May 2003 to change the data subsystem (DSS) from side A to side B. Following the switch to DSS-B ASAR operations resumed on 2 June 2003 with nominal data quality.
ASAR Antenna Reset #1 (September 2005)
A drift in the gain and phase of most Transmit and Receive Modules (TRMs) was seen from the start of the mission. Due to the accumulated gain and phase drift for a large percentage of the TRMs, the first of two antenna resets was performed on the 14th and 16th September 2005 in order to correct, as far as possible, the TRM drift.
ASAR Wide-Swath Burst Synchronisation (September 2006)
To perform interferometry with ASAR WS imagery, it is necessary for there to be an overlap in the bursts (in the azimuth direction) in the pair of products in each of the sub-swaths, otherwise it is not possible to generate an interferogram (since there will be no correlation between the pair of products as the same point on the ground would be imaged by a different part of the azimuth beam). The acquisition start times for ASAR WSM imagery were altered in September 2006 so that it started at discrete points around the orbit to maximise the chances of burst overlap between pairs of products. The revised planning strategy became operational from 17 September 2006 (orbit number 23783). Rosich et al. (2007) provide details of the Burst Synchronisation optimisation.
Recalibration of IM and AP modes (January 2007)
Following the successful re-deployment of three ASAR transponder from The Netherlands to Kalimantan (Indonesia), Resolute (Canada) and Ottawa (Canada) since mid-2006, many more transponder measurements have been made. This made it possible to perform a detailed analyses of the ASAR transponder relative radar cross-section (rcs) as a function of product type, swath and polarisation. This analysis showed the necessity of performing a re-calibration of IM and AP products via the generation of revised calibration constants. The radiometric re-calibration of ASAR Image and Alternating Polarisation modes was performed in December 2006 and January 2007 and is applicable for all products acquired since the start of the ASAR mission.
ASAR AP Swath Modifications (May 2009)
Unplanned shut-downs of the ASAR instrument had been occurring when acquiring alternating polarisation (AP) data since launch. Unplanned shut-downs are undesirable due to the possible degradation of the instrument hardware. The number of AP acquisitions was reduced significantly in early 2007 while the usage of instrument swath 5 (IS5) was suspended at the end of 2006, both to reduce the occurrence of these unplanned shut-downs. Changes were made to the ENVISAT ASAR AP swath characteristics at the end of May 2009 to reduce the number of shut-downs, thus enabling the resumption of AP IS5 data acquisitions and the increased usage of all other AP swaths. Details of the swath modifications can be found here.
ASAR Antenna Reset #2 (March 2010)
The decision to perform a second ASAR antenna reset activity during early 2010 was based on the gain and phase drifts that has occurred over the previous 4+ years for each of the 320 antenna TRMs (i.e. to bring the gain and phases close to pre-launch values). The upload of the files containing the required corrections took place as planned starting at 09:44 UT on 11 March 2010 during orbits 41974, 41975 and 41976, with a further upload on 17 March 2010 at 12:04:10 UT. Improvements were seen in all gains and phases after the reset, with values being relative to the start of the mission.
Envisat Orbit Change (October 2010)
In order the extend the ENVISAT mission beyond its nominal lifetime of 5 years and due to limited fuel remaining on-board (required for orbit manoeuvres and attitude control), it was decided to reduce the orbit altitude and cease some of the inclination orbit manoeuvres. The reduction in altitude was also the first step in placing ENVISAT in its end of life orbit. From 22 October 2010 the altitude of the orbit was reduced by about 17.4 km to 782.39 km. The Orbit Lowering Manoeuvre was completed on 2 Nov 2010 and the mission extension phase was entered.
Tandem missions, joint interferometric modes with ERS-2 and ENVISAT SAR instruments, occurred four times throughout ENVISAT's lifetime.
The first ENVISAT/ERS-2 SAR tandem mission took place from September 2007 to 14 February 14 2008.
The second ENVISAT/ERS-2 SAR tandem mission started on 23 November 2008 and ended in late April 2009 (both spacecraft acquired data over the same area just 28 minutes apart). The second campaign was due to end in late January 2009, but was extended through to late April 2009 due to the excellent performance of the satellites and the strong teamwork between the mission data teams at ESRIN and ESOC.
The third tandem campaign occurred from 14 February 2010 to 26 April 2010.
The fourth tandem campaign, from 6 July 2010 to 22 October 2010, focused on low-lying coastal areas, such as New Orleans in USA and the Po River Mouth in Italy.
Leap Frog Mode
From 26 November 2010 to 9 December 2010 an experimental Leap Frog mode was used, whereby the swath used from one wave mode imagette to another was switched between IS2 and IS4. This was repeated from 25 May 2011 to 6 June 2011, again alternating from IS2 and IS4. The polarisation remained constant, set to HH in a first step (2 weeks) and will move to VV for the remaining 2 weeks.